OpenStudy (anonymous):

Write an algebraic expression for the cost if n apples at .10 apiece and n+2 oranges at $0.15 apiece. @Roc07or

OpenStudy (anonymous):

if you want to ask someone for help just say there name with a @ in front of they're naem as in @bae7 or @roc07or

OpenStudy (anonymous):

*name

OpenStudy (aum):

1 apple costs $0.10 n apples will cost: $0.10 * n = 0.1n 1 orange costs $0.15 (n+2) oranges will cost: $0.15 * (n+2) = 0.15n + 0.3 Total cost = 0.1n + 0.15n + 0.3 = ?

OpenStudy (anonymous):

@aum thank you

OpenStudy (aum):

You are welcome.